OpenStudy (anonymous):

y = 2x + 3 and y = -x - 4

OpenStudy (anonymous):

2x + 3 = -x - 4 3x = -7 x=-7/3

OpenStudy (anonymous):

On a system of equations, don't forget to solve for y as well. The answer should be an ordered pair (x,y).